OpenStudy (18jonea):

Solve the following expression for x: 81=3(x+y)^3 x = 243 - y x = 27 - y x = 3 - y x = 9 - y

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

@Miracrown

OpenStudy (anonymous):

Ill try to help you but I'm not gonna put a answer if i think I'm wrong

OpenStudy (mathstudent55):

\(81=3(x+y)^3\) Divide both sides by 3. What do you get?

OpenStudy (18jonea):

27

OpenStudy (18jonea):

what three am i taking away on the side though the exponent or the first 3

OpenStudy (18jonea):

@mathstudent55

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (michele_laino):

hint: we can write this: \[\Large {\left( {x + y} \right)^3} = \frac{{81}}{3} = 27\]

OpenStudy (18jonea):

ok then how do i get rid of the exponent

OpenStudy (michele_laino):

you have to take the third rooth of both sides

OpenStudy (18jonea):

so how would i do that?

OpenStudy (18jonea):

how do you takw the 3 root of both sides

OpenStudy (mathstudent55):

Correct. \(27=(x+y)^3\) Now we notice that \(27 = 3^3\), so we write: \(3^3 = (x + y)^3\) If the third power of 3 equals the third power of x + y, then 3 must equal x + y. You can also take the cubic root of both sides. \(3 = x + y\) There is only one step left. Subtract y from both sides.

OpenStudy (18jonea):

so would it be 3-y=x @Michele_Laino or 9-y= x

OpenStudy (michele_laino):

it is the first one, namely x+y=3

OpenStudy (18jonea):

ok

OpenStudy (18jonea):

The following equation is used to determine the kinetic energy of an object: E=1/2mv^2. Solve for m:

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

would it be 2E/ v^2

OpenStudy (18jonea):

@mathstudent55

OpenStudy (michele_laino):

correct!

OpenStudy (18jonea):

Thank you

OpenStudy (18jonea):

Write an expression for the following description: 23 less than the product of x and 6. 6x-23 is that right

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (michele_laino):

yes!